Jump to main content.


Drake Chemical

EPA Daily Situation Reports for July 1998

U. S. ARMY CORPS OF ENGINEERS
DRAKE CHEMICAL SUPERFUND SITE

182 MYRTLE STREET
LOCK HAVEN, PA 17745
PHONE 717-748-4505
FAX 717-748-7549

Friday, July 31, 1998

FOR DATES: From 7:00 a.m. Thursday, July 30, 1998 to 7:00 a.m. Friday, July 31, 1998.

WEATHER: Clear and mild. The temperature range was 58oF to 81oF.

I. CURRENT SITUATION: The Thermal Destruction Facility (TDF) is currently off-line to allow for the replacement of the kiln feed screws, and perform maintenance on the Induced Draft (ID) fan and other components. The contractor continues to work 24 hours a day, performing maintenance on the TDF equipment. The contractor is performing related activities such as soil feed preparation and debris separation.

II. ACTIONS TAKEN: Continued TDF maintenance. Continued feed preparation, WWTP operations, removal of the top 2 feet of perimeter sheetpile along the railroad property, lime mixing, feed preparation, debris separation, dust control, rock and inorganic debris cleaning, and rock backfilling at the bottom of the excavation. Continued real time perimeter air monitoring activities. Received bottom ash analytical report (bot-ash-07-26-98-178) and fly ash report (fly-ash-07-25-98-032). The chemical analysis indicated that the bottom ash met project specifications for unrestricted backfilling on-site. The chemical analyses for the fly ash indicated that the fly ash materials met the project requirements for restricted backfilling on-site, due to the concentrations of chromium and selenium being greater than the PADEP drinking water standard, but less than 25 times the standard. The fly ash will be restricted to placement above elevation 553.

III. FUTURE PLANS: Major activities for the next week include the restart and operation of the TDF, excavation, drum handling, feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, rock cleaning, rock backfilling, and inorganic debris decontamination.

IV. PROBLEMS NOTED OR ANTICIPATED: None.

V. NOTED VISITORS: Sandy Downs (PADEP)

PREPARED BY: Steve Gibson for Dave Modricker (USACE)


Thursday, July 30, 1998

FOR DATES: From 7:00 a.m. Wednesday, July 29, 1998 to 7:00 a.m. Thursday, July 30, 1998.

WEATHER: Partly Cloudy with afternoon rain. The temperature range was 56oF to 84oF.

I. CURRENT SITUATION: The Thermal Destruction Facility (TDF) is currently off-line to allow for the replacement of the kiln feed screws, and perform maintenance on the Induced Draft (ID) fan and other components. The contractor continues to work 24 hours a day, performing maintenance on the TDF equipment. The contractor is performing related activities such as soil feed preparation and debris separation.

II. ACTIONS TAKEN: Continued TDF maintenance. Continued WWTP operations, removal of the top 2 feet of perimeter sheetpile along the railroad property, lime mixing, feed preparation, debris separation, dust control, drum bulking during the night-shift, rock and inorganic debris cleaning, and rock backfilling. Continued real time perimeter air monitoring activities.

III. FUTURE PLANS: Major activities for the next week include the restart and operation of the TDF, excavation, drum handling, drum bulking, feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, rock cleaning, rock backfilling, and inorganic debris decontamination.

IV. PROBLEMS NOTED OR ANTICIPATED: None.

V. NOTED VISITORS: None.

PREPARED BY: Steve Gibson for Dave Modricker (USACE)


Wednesday, July 29, 1998

FOR DATES: From 7:00 a.m. Tuesday, July 28, 1998 to 7:00 a.m. Wednesday, July 29, 1998.

WEATHER: Hazy, partly cloudy. The temperature range was 63oF to 83oF.

I. CURRENT SITUATION: The Thermal Destruction Facility (TDF) is currently off-line to allow for the replacement of the kiln feed screws, and perform maintenance on the Induced Draft (ID) fan and other components. The contractor continues to work 24 hours a day, performing maintenance on the TDF equipment. The contractor is performing related activities such as soil feed preparation and debris separation.

II. ACTIONS TAKEN: Continued TDF maintenance. Continued WWTP operations, removal of the top 2 feet of perimeter sheetpile along the railroad property, lime mixing, feed preparation, debris separation, dust control, drum bulking during the night-shift, rock and inorganic debris cleaning, and rock backfilling. Continued real time perimeter air monitoring activities. Received bottom ash analytical reports for bot-ash-07-24-98-176. The chemical analysis indicated that the thermally treated materials met project specifications for unrestricted backfilling on-site (i.e.: above elevation 547).

III. FUTURE PLANS: Major activities for the next week include the restart and operation of the TDF, excavation, drum handling, drum bulking, feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, rock cleaning, rock backfilling, and inorganic debris decontamination.

IV. PROBLEMS NOTED OR ANTICIPATED: None.

V. NOTED VISITORS: Mike Welch and Sandy Downs (PADEP), Kevin Fanning (CCDES)

PREPARED BY: Steve Gibson for Dave Modricker (USACE)


Tuesday, July 28, 1998

FOR DATES: From 7:00 a.m. Monday, July 27, 1998 to 7:00 a.m. Tuesday, July 28, 1998.

WEATHER: Mostly clear. The temperature range was 58oF to 82oF.

I. CURRENT SITUATION: The Thermal Destruction Facility (TDF) is currently off-line to allow for the replacement of the kiln feed screws, and perform maintenance to the Induced Draft (ID) fan and other components. The contractor continues to work 24 hours a day, performing maintenance on the TDF equipment. The contractor is performing related activities such as soil feed preparation and debris separation.

II. ACTIONS TAKEN: Continued TDF maintenance. The Thermal Relief Valve (TRV) was opened at 0740 hours. OHM's Quality Control staff made the proper notifications. Continued WWTP operations, lime mixing, feed preparation, debris separation, dust control, rock and inorganic debris cleaning, and rock backfilling. Drum bulking activities were restarted on this date. Continued real time perimeter air monitoring activities.

Received bottom ash and fly ash analytical reports for bot-ash-07-19-98-171, bot-ash-07-21-98-173, bot-ash-07-22-98-174, bot-ash-07-23-98-175, and fly-ash-07-20-98-031. The chemical analyses indicated that the materials sampled met the project requirements for unrestricted backfilling on-site, except for bot-ash-07-22-98-174 and fly-ash-07-20-98-031. Bottom ash sample 07-22-98-174 met the project requirements for restricted backfilling (i.e. greater than elevation 553) due to the concentrations of lead greater than the PADEP drinking water standard, but less than 25 times the standard. Fly ash sample 07-20-98-031 met the project requirements for restricted backfilling (i.e. greater than elevation 553) due to the concentrations of chromium, cadmium, and selenium greater than the PADEP drinking water standard, but less than 25 times the standard.

III. FUTURE PLANS: Major activities for the next week include the restart and operation of the TDF, excavation, drum handling, drum bulking, feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, rock cleaning, rock backfilling, and inorganic debris decontamination.

IV. PROBLEMS NOTED OR ANTICIPATED: None.

V. NOTED VISITORS: MG Jerry Sinn, MJ Greg Stinner, Joseph D'Agosta, Lloyd Caldwell, K. Charlie Chung, Mohan Singh (USACE), Mike Welch, Robert Yowell (PADEP), Enzo Zoratto, Glenn Phelps, Matt Radek (OHM)

PREPARED BY: Steve Gibson for Dave Modricker (USACE)


Monday, July 27, 1998

FOR DATES: From 7:00 a.m. Sunday, July 26, 1998 to 7:00 a.m. Monday, July 27, 1998.

WEATHER: Mostly clear. The temperature range was 57oF to 79oF.

I. CURRENT SITUATION: The Thermal Destruction Facility (TDF) is currently off-line to allow for maintenance. Feed was stopped at 1836 hours in preparation for the scheduled maintenance. The TDF was on-line for 11.5 hours and 372.9 tons of contaminated soil was thermally treated during the time frame of this report. Site soil was processed at an average rate of 32.4 tons per hour. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: Continued TDF operations, prior to shutdown. Minor excavation, lime mixing, feed preparation, debris separation, dust control, rock and inorganic debris cleaning, ash handling, and ash backfilling are on-going activities. Continued real time perimeter air monitoring activities.

III. FUTURE PLANS: Major activities for the next week include maintenance of the TDF, excavation, drum handling, drum bulking, feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, rock cleaning, and inorganic debris decontamination.

The scheduled TDF cold shutdown started at 1836 hours. The shutdown is planned to last approximately one week. While the unit is shut down, preventative maintenance will be performed on the Induced Draft (ID) fan, the feed screws will be replaced, and general maintenance on the TDF will be performed. A scheduled Thermal Relief Valve (TRV) opening occurred at 2130 hours during the interlock checks. An extended TRV opening is planned for Monday, once maintenance activities begin.

IV. PROBLEMS NOTED OR ANTICIPATED: None.

V. NOTED VISITORS: None.

PREPARED BY: Danny Zion for Dave Modricker (USACE)


Sunday, July 26, 1998

FOR DATES: From 7:00 a.m. Saturday, July 25, 1998 to 7:00 a.m. Sunday, July 26, 1998.

WEATHER: Mostly clear. The temperature range was 56oF to 77oF.

I. CURRENT SITUATION: The contractor is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 22.5 hours and 729.7 tons of contaminated soil was thermally treated during the time frame of this report. Off-line time is attributed to a broken drive chain on the feed screws, which accumulated 1 hour and 24 minutes of down time. Site soil was processed at an average rate of 32.4 tons per hour. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: Continued TDF operations, minor excavation, lime mixing, feed preparation, debris separation, dust control, rock and inorganic debris cleaning, ash handling, and ash backfilling. Continued real time perimeter air monitoring activities and began the 24-hour time weighted air monitoring at the off-site stations.

III. FUTURE PLANS: Major activities for the next week include maintenance of the TDF; excavation, drum handling, drum bulking, feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, rock cleaning, and inorganic debris decontamination.

There is a scheduled TDF cold shutdown planned to start this evening at approximately 7:00 p.m. The shutdown is planned to last approximately one week. While the unit is shut down, preventative maintenance will be performed on the Induced Draft (ID) fan, the feed screws will be replaced, and general maintenance on the TDF will be performed. A Thermal Relief Valve (TRV) opening is scheduled for Sunday evening during the interlock check at approximately 9:00 p.m., and an extended TRV opening is planned for Monday, once maintenance activities begin.

IV. PROBLEMS NOTED OR ANTICIPATED: None.

V. NOTED VISITORS: None.

PREPARED BY: Greg Beauseigneur for Dave Modricker (USACE)


Saturday, July 25, 1998

FOR DATES: From 7:00 a.m. Friday, July 24, 1998 to 7:00 a.m. Saturday, July 25, 1998.

WEATHER: Partly cloudy. The temperature range was 56oF to 75oF.

I. CURRENT SITUATION: The contractor is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 23.5 hours and 779.5 tons of contaminated soil was thermally treated during the time frame of this report. Off-line time is attributed to the feed screws becoming overloaded with soil materials. The majority of the stoppages lasted for 1 to 2 minutes each. Contaminated soil was processed at an average rate of 33 tons per hour. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: Continued TDF operations, minor excavation, lime mixing, feed preparation, debris separation, dust control, rock and inorganic debris cleaning, ash handling, and ash backfilling. Continued real time perimeter air monitoring activities. Received bottom ash analytical report (bot-ash-07-20-98-172) for bin 2. The chemical analyses indicated that the thermally treated soil met the project requirements for unrestricted backfilling on-site (ie: above elevation 547).

III. FUTURE PLANS: Major activities for the next week include maintenance of the TDF; excavation, drum handling, drum bulking, feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, rock cleaning, and inorganic debris decontamination.

There is a scheduled TDF cold shutdown planned to start during the evening hours of Sunday, July 26, 1998. The shutdown is planned to last approximately one week. While the unit is shut down, preventative maintenance will be performed on the Induced Draft (ID) fan, the feed screws will be replaced, and general maintenance on the TDF will be performed. A Thermal Relief Valve (TRV) opening is scheduled for Sunday evening during the interlock check, and an extended TRV opening is planned for Monday, once maintenance activities begin.

IV. PROBLEMS NOTED OR ANTICIPATED: None.

V. NOTED VISITORS: None.

PREPARED BY: Steve Gibson for Dave Modricker (USACE)


Friday, July 24, 1998

FOR DATES: From 7:00 a.m. Thursday, July 23, 1998 to 7:00 a.m. Friday, July 24, 1998.

WEATHER: Overcasts with periods of rain. The temperature range was 63oF to 85oF.

I. CURRENT SITUATION: The contractor is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 21.6 hours and 727.6 tons of contaminated soil was thermally treated during the time frame of this report. Down time is attributed to the repairing of the weighbelt, which was torn by a piece of steel debris. Site soil was processed at an average rate of 33.7 tons per hour. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: Continued TDF operations, WWTP operations, excavation, lime mixing, feed preparation, debris separation, dust control, rock and inorganic debris cleaning, rock backfilling, ash handling, and ash backfilling. Continued real time perimeter air monitoring activities.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF; continuation of excavation, drum handling, drum bulking, feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination.

There is a scheduled TDF cold shutdown planned to start during the evening hours of Sunday, July 26, 1998. The shutdown is planned to last approximately one week. While the unit is shut down, preventative maintenance is to be performed on the induced draft (ID) fan, the feed screws are to be replaced, and general maintenance on the TDF is to be performed. A Thermal Relief Valve (TRV) opening is scheduled for Sunday evening during the interlock check, and an extended TRV opening is planned for Monday, once maintenance activities begin.

IV. PROBLEMS NOTED OR ANTICIPATED: None.

V. NOTED VISITORS: Sandy Downs (PADEP)

PREPARED BY: Steve Gibson for Frank Cleaver (USACE)


Thursday, July 23, 1998

FOR DATES: From 7:00 a.m. Wednesday, July 22, 1998 to 7:00 a.m. Thursday, July 23, 1998.

WEATHER: Clear. The temperature range was 69oF to 84oF.

I. CURRENT SITUATION: The contractor is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 23.9 hours and 814.7 tons of contaminated soil were thermally treated during the time frame of this report. Down time is attributed to feed belt calibration check. Site soil was processed at an average rate of 34 tons per hour. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: Continued TDF operations, WWTP operations, excavation, lime mixing, feed preparation, debris separation, dust control, rock and inorganic debris cleaning, rock backfilling, ash handling, and ash backfilling. Continued real time perimeter air monitoring activities. Received bottom ash analytical reports for bot-ash-07-16-98-168, bot-ash-07-17-98-169, and bot-ash-07-18-98-170. Bottom ash tested met project specifications for backfill.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF; continuation of excavation, drum handling, drum bulking, feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination. There is a scheduled TDF cold shutdown planned to start Sunday, July 26, 1998 and last approximately one week. While the unit is shut down, preventative maintenance is to be performed on the induced draft (ID) fan, the feed screws are to be replaced and general maintenance on the TDF is to be performed.

IV. PROBLEMS NOTED OR ANTICIPATED: None.

V. NOTED VISITORS: None.

PREPARED BY: Danny Zion for Frank Cleaver (USACE)


Wednesday, July 22, 1998

FOR DATES: From 7:00 a.m. Tuesday, July 21, 1998 to 7:00 a.m. Wednesday, July 22, 1998.

WEATHER: Humid with an afternoon thunderstorm. The temperature range was 70oF to 83oF.

I. CURRENT SITUATION: The contractor is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 22.9 hours and 785.6 tons of contaminated soil were thermally treated during the time frame of this report. Down time is attributed to local lightning storm activity. Site soil was processed at an average rate of 34.3 tons per hour. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: Continued TDF operations, WWTP operations, lime mixing, feed preparation, debris separation, dust control, rock and inorganic debris cleaning, rock backfilling, ash handling, and ash backfilling. Continued real time perimeter air monitoring activities. Received fly ash analytical report for fly-ash-07-15-98-030. The chemical analysis indicated that the thermally treated soil met the project requirements for restricted backfilling on-site. Due to the concentrations of Chromium and Selenium being greater than the PADEP drinking water standard, but less than 25 times the standard, the fly ash is restricted to backfill above elevation 553.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF; continuation of excavation, drum handling, drum bulking, feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination.

IV. PROBLEMS NOTED OR ANTICIPATED: None.

V. NOTED VISITORS: Attendees of the weekly progress meeting.

PREPARED BY: Steve Gibson for Bill Werntges (USACE)


Tuesday, July 21, 1998

FOR DATES: From 7:00 a.m. Monday, July 20, 1998 to 7:00 a.m. Tuesday, July 21, 1998.

WEATHER: Humid and partly cloudy. The temperature range was 67oF to 84oF.

I. CURRENT SITUATION: The contractor is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 23.8 hours and 803 tons of contaminated soil were thermally treated during the time frame of this report. Site soil was processed at an average rate of 37.1 tons per hour. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: Continued TDF operations, WWTP operations, lime mixing, feed preparation, debris separation, dust control, ash handling, and ash backfilling. Excavation activities were restarted on this date. Continued real time perimeter air monitoring activities and completed the 24-hour sampling at the off-site air monitoring stations. Received bottom ash analytical report for bot-ash-07-15-98-167. The chemical analysis indicated that the thermally treated soil met the project requirements for unrestricted backfilling on-site.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF; continuation of excavation, drum handling, drum bulking, feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock washing & inorganic debris decontamination.

IV. PROBLEMS NOTED OR ANTICIPATED: None.

V. NOTED VISITORS: None.

PREPARED BY: Steve Gibson for Bill Werntges (USACE)


Monday, July 20, 1998

FOR DATES: From 7:00 a.m. Sunday, July 19, 1998 to 7:00 a.m. Monday, July 20, 1998.

WEATHER: Partly sunny. The temperature range was 56oF to 83oF.

I. CURRENT SITUATION: The contractor is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 21 hours, 27 minutes and 803 tons of contaminated soils were thermally treated during the time frame of this report. Site soil was processed at an average rate of 37.4 tons per hour. The majority of the downtime was due to thunderstorms in the area. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: Continued TDF operations, lime mixing, feed preparation, debris separation, dust control, ash handling, and ash backfilling. Continued real time perimeter air monitoring activities.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF; continuation of excavation, drum handling, drum bulking, feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock washing & inorganic debris decontamination.

IV. PROBLEMS NOTED OR ANTICIPATED: None.

V. NOTED VISITORS: Site tours were given from about 10 a.m. to 2 p.m. for employee's relatives and friends.

PREPARED BY: Danny Zion for Bill Werntges (USACE)


Sunday, July 19, 1998

FOR DATES: From 7:00 a.m. Saturday, July 18, 1998 to 7:00 a.m. Sunday, July 19, 1998.

WEATHER: Clear and mild. The temperature range was 57.2oF to 82.2oF.

I. CURRENT SITUATION: The contractor is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 23 hours, 32 minutes and 878.5 tons of contaminated soil were thermally treated during the time frame of this report. Site soil was processed at an average rate of 37.33 tons per hour. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: Continued TDF operations, lime mixing, feed preparation, debris separation, dust control, ash handling, and ash backfilling. Continued real time perimeter air monitoring activities.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF; continuation of excavation, drum handling, drum bulking, feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock washing & inorganic debris decontamination.

IV. PROBLEMS NOTED OR ANTICIPATED: None.

V. NOTED VISITORS: None

PREPARED BY: Richard Lupold for David Modricker (USACE)


Saturday, July 18, 1998

FOR DATES: From 7:00 a.m. Friday, July 17, 1998 to 7:00 a.m. Saturday, July 18, 1998.

WEATHER: Hazy with thunderstorms. The temperature range was 62oF to 83oF.

I. CURRENT SITUATION: The contractor is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 22.3 hours and 866.5 tons of contaminated soils were thermally treated during the time frame of this report. Off-line time is attributed to the manual shutdown of feed during two local lightning storm events. Site soil was processed at an average rate of 38.9 tons per hour. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: Continued TDF operations, lime mixing, feed preparation, debris separation, dust control, ash handling, and ash backfilling. Continued real time perimeter air monitoring activities. Debris screening resumed on this date, once trommel screener #2 had been repaired. Received bottom ash analytical report for bot-ash-07-13-98-165. The chemical analyses indicated that the thermally treated soil met all requirements for unrestricted backfilling on-site (ie: above elevation 547).

III. FUTURE PLANS: Major activities for the next week include operation of the TDF; continuation of excavation, drum handling, drum bulking, feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock washing & inorganic debris decontamination.

IV. PROBLEMS NOTED OR ANTICIPATED: None.

V. NOTED VISITORS: Mike Welch and Sandy Downs (PADEP)

PREPARED BY: Steve Gibson for David Modricker (USACE)


Friday, July 17, 1998

FOR DATES: From 7:00 a.m. Thursday, July 16, 1998 to 7:00 a.m. Friday, July 17, 1998.

WEATHER: Hazy and partly cloudy. The temperature range was 68oF to 82oF.

I. CURRENT SITUATION: The contractor is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 23 hours 46 minutes and 952.4 tons of contaminated soil was thermally treated during the time frame of this report. Site soil was processed at an average rate of 40.1 tons per hour. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: The team of inspectors, led by OSHA, concluded the site audit of the health and safety program and work practices. The inspection was part of a national program developed by USEPA and OSHA to inspect hazardous waste Superfund sites utilizing incineration. Their findings will be forwarded in a report to USEPA Region III. Continued TDF operations, feed preparation, debris separation, dust control, ash handling, and ash backfilling. Continued real time perimeter air monitoring activities. New excavation of contaminated soil from the south side of the stockpile was not accomplished during the time frame of this report. The trommel screeners in the debris handling building are being repaired. While the trommels are out of use, screening of feed is being done with a power screen in the feed handling building.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF; continuation of excavation, drum handling, drum bulking, feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock washing & inorganic debris decontamination.

IV. PROBLEMS NOTED OR ANTICIPATED: None.

V. NOTED VISITORS: Earl Cook (OSHA), Cynthia Duffield and Nina Baird (ATL), Richard Wright (USACE Washington Headquarters), Charles Bragdon (USACE Baltimore District), David Mummert (OHM CIH), Sandy Downs (PADEP)

PREPARED BY: David Modricker (USACE)


Thursday, July 16, 1998

FOR DATES: From 7:00 a.m. Wednesday, July 15, 1998 to 7:00 a.m. Thursday, July 16, 1998.

WEATHER: cloudy. The temperature range was 66oF to 80oF.

I. CURRENT SITUATION: The contractor is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 23 hours 56 minutes and 959.2 tons of contaminated soil was thermally treated during the time frame of this report. Site soil was processed at an average rate of 40 tons per hour. Related ongoing site activities include feed preparation and debris separation. A team of inspectors, led by OSHA, is on site performing an audit of the health and safety program and work practices. The inspection is part of a national program developed by USEPA and OSHA to inspect hazardous waste Superfund sites utilizing incineration.

II. ACTIONS TAKEN: Continued TDF operations, feed preparation, debris separation, dust control, ash handling, and ash backfilling. Continued real time perimeter air monitoring activities. New excavation of contaminated soil from the south side of the stockpile was not accomplished during the time frame of this report. The trommel screeners in the debris handling building are being repaired. While the trommels are out of use, screening of feed is being done with a power screen in the feed handling building.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF; continuation of excavation, drum handling, drum bulking, feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock washing & inorganic debris decontamination.

IV. PROBLEMS NOTED OR ANTICIPATED: None.

V. NOTED VISITORS: Earl Cook (OSHA), Cynthia Duffield and Nina Baird (ATL), Richard Wright (USACE Washington Headquarters), Charles Bragdon (USACE Baltimore District), David Mummert (OHM CIH), Kevin Fanning and Rick Scheesley of Clinton County DES, and Chris McBride (Focus Engineering).

PREPARED BY: Danny Zion for Dave Modricker (USACE)


Wenesday, July 15, 1998

FOR DATES: From 7:00 a.m. Tuesday, July 14, 1998 to 7:00 a.m. Wednesday, July 15, 1998.

WEATHER: cloudy. The temperature range was 62oF to 84oF.

I. CURRENT SITUATION: The contractor is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 23 hours 46 minutes and 953.5 tons of contaminated soil were thermally treated during the time frame of this report. Site soil was processed at an average rate of 40 tons per hour. Related ongoing site activities include feed preparation and debris separation. A team of inspectors, led by OSHA, is on site performing an audit of the health and safety program and work practices. The inspection is part of a national program developed by USEPA and OSHA to inspect hazardous waste Superfund sites utilizing incineration.

II. ACTIONS TAKEN: Continued TDF operations, feed preparation, debris separation, dust control, ash handling, and ash backfilling. Continued real time perimeter air monitoring activities. New excavation of contaminated soil from the south side of the stockpile was not accomplished during the time frame of this report. One of the trommel screeners in the debris handling building is being repaired and a grizzly is being installed on the second trommel. While the trommels are out of use, screening of feed is being done with a power screen in the feed handling building. All three runs of stack testing for dioxins and furans were completed on July 14, 1998 by subcontractor Air-Recon.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF; continuation of excavation, drum handling, drum bulking, feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock washing & inorganic debris decontamination.

IV. PROBLEMS NOTED OR ANTICIPATED: None

V. NOTED VISITORS: S. Downs, L. Newcomer, S. Webster, and M. Welch of PADEP. Earl Cook (OSHA), Cynthia Duffield and Nina Baird (ATL), Richard Wright (USACE Washington Headquarters), Charles Bragdon (USACE Baltimore District), Hugh Kaufman (USEPA), Marc Laquercia (Booz-Allen), David Mummert (OHM CIH), and Chris McBride (Focus Engineering).

PREPARED BY: Danny Zion for Dave Modricker (USACE)


Tuesday, July 14, 1998

FOR DATES: From 7:00 a.m. Monday, July 13, 1998 to 7:00 a.m. Tuesday, July 14, 1998.

WEATHER: Sunny. The temperature range was 56oF to 84oF.

I. CURRENT SITUATION: The contractor is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 23 hours 43 minutes and 951 tons of contaminated soil were thermally treated during the time frame of this report. Site soil was processed at an average rate of 40 tons per hour. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: Continued TDF operations, feed preparation, debris separation, dust control, ash handling, and ash backfilling. Continued real time perimeter air monitoring activities. Received bottom ash analytical test results for bot-ash-07-08-98-160, bot-ash-07-09-98-161A, and bot-ash-07-09-98-161B. Bottom ash met project specifications for backfill at or above elevation 547. New excavation of contaminated soil from the south side of the stockpile was not accomplished during the time frame of this report. Trommel screeners in the debris handling building are under repair. While the trommels are down, screening of feed is being done with a power screen in the feed handling building.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF; continuation of excavation, drum handling, drum bulking, feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock washing & inorganic debris decontamination. Stack testing for dioxins and furans began on July 14, 1998.

IV. PROBLEMS NOTED OR ANTICIPATED: Short term and long term plans to handle debris on-site are being evaluated.

V. NOTED VISITORS: Earl Cook (OSHA), Cynthia Duffield and Nina Baird (ATL), Richard Wright (USACE Washington Headquarters), Charles Bragdon (USACE Baltimore District), David Mummert (OHM CIH), and Chris McBride (Focus Engineering).

PREPARED BY: Danny Zion for Dave Modricker (USACE)


Monday, July 13, 1998

FOR DATES: From 7:00 a.m. Sunday, July 12, 1998 to 7:00 a.m. Monday, July 13, 1998.

WEATHER: Sunny. The temperature range was 50oF to 78oF.

I. CURRENT SITUATION: The contractor is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 23 hours 49 minutes and 954.8 tons of contaminated soil was thermally treated during the time frame of this report. Site soil was processed at an average rate of 40 tons per hour. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: Continued TDF operations, feed preparation, debris separation, dust control, ash handling, and ash backfilling. Continued real time perimeter air monitoring activities. Received bottom ash analytical test results for bot-ash-07-07-98-159. Bottom ash met project specifications for backfill at or above elevation 547.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF; continuation of excavation, drum handling, drum bulking, feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock washing & inorganic debris decontamination. Stack testing for dioxins and furans is scheduled for July 14 and 15, 1998.

IV. PROBLEMS NOTED OR ANTICIPATED: None

V. NOTED VISITORS: None

PREPARED BY: Danny Zion for Dave Modricker (USACE)


Sunday, July 12, 1998

FOR DATES: From 7:00 a.m. Saturday, July 11, 1998 to 7:00 a.m. Sunday, July 12, 1998.

WEATHER: Sunny and mild. The temperature range was 50.3oF to 78.2oF.

I. CURRENT SITUATION: The contractor is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 23 hours 56 mins and 959.5 tons of contaminated soil were thermally treated during the time frame of this report. Site soil was processed at an average rate of 40 tons per hour. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: Continued TDF operations, feed preparation, debris separation, dust control, ash handling, and ash backfilling. Continued real time perimeter air monitoring activities. Received bottom ash analytical test results for bot-ash-07-06-98-158 and fly ash analytical test results for fly-ash-07-04-98-027. Bottom ash met project specifications for backfill at or above elevation 547 and the fly ash met project specifications for backfill at or above elevation 553.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF; continuation of excavation, drum handling, drum bulking, feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock washing & inorganic debris decontamination. Stack testing for dioxins and furans is scheduled for July 14 and 15, 1998.

IV. PROBLEMS NOTED OR ANTICIPATED: None

V. NOTED VISITORS: None

PREPARED BY: Richard Lupold for Dave Modricker (USACE)


Saturday, July 11, 1998

FOR DATES: From 7:00 a.m. Friday, July 10, 1998 to 7:00 a.m. Saturday, July 11, 1998.

WEATHER: Sunny with few clouds. The temperature range was 50oF to 77oF.

I. CURRENT SITUATION: The contractor is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 23.8 hours and 943.3 tons of contaminated soil were thermally treated during the time frame of this report. Site soil was being processed at an average rate of 39.6 tons per hour. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: Continued TDF operations, WWTP operations, feed preparation, debris separation, oversized rock-concrete debris washing, ash handling, and ash backfilling. The change out of the carbon for the WWTP carbon vessels was completed Friday morning. Continued real time perimeter air monitoring activities. Received bottom ash analytical test results for bot-ash-07-03-98-155. Bottom ash met project specifications for backfill at or above elevation 547.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF; continuation of excavation, drum handling, drum bulking, feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock washing & inorganic debris decontamination. Stack testing for dioxins and furans is scheduled for July 14 and 15, 1998.

IV. PROBLEMS NOTED OR ANTICIPATED: None

V. NOTED VISITORS: Sandy Downs (PADEP)

PREPARED BY: Steve Gibson for Dave Modricker (USACE)


Friday, July 10, 1998

FOR DATES: From 7:00 a.m. Thursday, July 9, 1998 to 7:00 a.m. Friday, July 10, 1998.

WEATHER: Overcast, clearing in the afternoon. The temperature range was 63oF to 80oF.

I. CURRENT SITUATION: The contractor is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 22 hours and 864.3 tons of contaminated soil were thermally treated during the time frame of this report. Site soil was being processed at an average rate of 39.2 tons per hour (tph). The contractor stopped feed at 1411 hours. Feed resumed at 1601 hours. The feed rate was reduced from approximately 42 tph to 38 tph.

II. ACTIONS TAKEN: Continued TDF operations, WWTP operations, feed preparation, debris separation, large rock and concrete debris washing & backfilling, ash handling, and ash backfilling. Continued real time perimeter air monitoring activities. The carbon in the carbon vessels at the WWTP was changed on this date. Received bottom ash analytical reports (bot-ash-7-4-98-156 and bot-ash-7-5-98-157) for bins 6 and 7. The analytical reports for both samples indicated levels of BNA greater than the allowable backfill standard. Ash from both bins will be returned to the exclusion zone for further thermal treatment. As noted above, the TDF feed rate was reduced to 38 tph and is expected to remain between 38 tph and 40 tph while OHM reviews parameters that affect ash quality.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF; continuation of excavation, drum handling, drum bulking, feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock washing & inorganic debris decontamination. Stack testing for dioxins and furans is scheduled for July 14 and 15, 1998.

IV. PROBLEMS NOTED OR ANTICIPATED: Chemical quality control issues with the analysis of Fenac in the ash by the contractor's primary lab was a concern on this date. As of the issuance of this report, the chemical quality control issue with the analysis of Fenac is resolved.

V. NOTED VISITORS: None

PREPARED BY: Steve Gibson and Dave Modricker (USACE)


Thursday, July 9, 1998

FOR DATES: From 7:00 a.m. Wednesday, July 8, 1998 to 7:00 a.m. Thursday, July 9, 1998.

WEATHER: Overcast with periods of heavy rain. The temperature range was 63oF to 68oF.

I. CURRENT SITUATION: The contractor is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 23.9 hours and 1009.4 tons of contaminated soil were thermally treated during the time frame of this report. Site soil was being processed at an average rate of 42.2 tons per hour. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: Continued TDF operations, WWTP operations, feed preparation, debris separation, ash handling, and ash backfilling. Continued real time perimeter air monitoring activities and concluded the 24-hour off-site air sampling. Minor erosion control earthwork was performed at the top and bottom of the excavation face.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF; continuation of excavation, drum handling, drum bulking, feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock washing & inorganic debris decontamination. Stack testing for dioxins and furans is scheduled for July 14 and 15, 1998.

IV. PROBLEMS NOTED OR ANTICIPATED: None

V. NOTED VISITORS: None

PREPARED BY: Steve Gibson and Dave Modricker (USACE)


Wednesday, July 8, 1998

FOR DATES: From 7:00 a.m. Tuesday, July 7, 1998 to 7:00 a.m. Wednesday, July 8, 1998.

WEATHER: Humid and sunny with evening rain. The temperature range was 62oF to 78oF.

I. CURRENT SITUATION: The contractor is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 23.9 hours and 1013.4 tons of contaminated soil was thermally treated during the time frame of this report. Off-line time is attributed to the performance of the daily weight belt calibration. Site soil was being processed at a rate of 42.4 tons per hour. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: Continued TDF operations, WWTP operations, excavation, feed preparation, debris separation, dust control, ash handling, and ash backfilling. Continued the placement odor suppressant foam on the excavation face. Continued real time perimeter air monitoring activities and began the 24-hour off-site air sampling.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF; continuation of excavation, drum handling, drum bulking, feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock washing & inorganic debris decontamination. Stack testing for dioxins and furans is scheduled for July 14 and 15, 1998.

IV. PROBLEMS NOTED OR ANTICIPATED: None

V. NOTED VISITORS: Bruce Rindell, Sandy Downs, Larry Newcomer, Steve Webster (PADEP), and Members of the City of Lock Haven Environmental Advisory Committee.

PREPARED BY: Steve Gibson and Dave Modricker (USACE)


Tuesday, July 7, 1998

FOR DATES: From 7:00 a.m. Monday, July 6, 1998 to 7:00 a.m. Tuesday, July 7, 1998.

WEATHER: Partly cloudy to sunny. The temperature range was 54oF to 79oF.

I. CURRENT SITUATION: The contractor is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 23.1 hours and 977.7 tons of contaminated soil was thermally treated during the time frame of this report. Off-line time is attributed to the performance of the daily weight belt calibration, the stoppage of the Dry Ash Drag conveyor belt, and low Secondary Combustion Chamber temperature. Site soil was being processed at a rate of 42.3 tons per hour. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: Continued TDF operations, WWTP operations, excavation, feed preparation, debris separation, dust control, ash handling, and ash backfilling. Continued the placement odor suppressant foam on the excavation face. Continued real time perimeter air monitoring and daily off-site air monitoring activities.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF; continuation of excavation, drum handling, drum bulking, feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock washing & inorganic debris decontamination. Stack testing for dioxins and furans is scheduled for July 14 and 15, 1998.

IV. PROBLEMS NOTED OR ANTICIPATED: None

V. NOTED VISITORS: Sandy Downs (PADEP)

PREPARED BY: Steve Gibson and Bill Werntges (USACE)


Monday, July 6, 1998

FOR DATES: From 7:00 a.m. Sunday, July 5, 1998 to 7:00 a.m. Monday, July 6, 1998.

WEATHER: Partly cloudy. The temperature range was 53oF to 78oF.

I. CURRENT SITUATION: The contractor is working 24 hours a day, 7 days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 23 hours 55 minutes and 1014 tons of contaminated soils were treated during the time frame of this report. Site soil was processed at an average rate of 42.4 tons per hour. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: Continued TDF operations, feed preparation, debris separation, dust control, ash handling, and ash backfilling. Continued with real time perimeter air monitoring and off-site air monitoring activities.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF; continuation of excavation, drum handling, drum bulking, feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock washing & inorganic debris decontamination. Stack testing for dioxins and furans is scheduled for July 14 and 15, 1998.

IV. PROBLEMS NOTED OR ANTICIPATED: None

V. NOTED VISITORS: None

PREPARED BY: Danny Zion for David Modricker (USACE).


Sunday, July 5, 1998

FOR DATES: From 7:00 a.m. Saturday, July 4, 1998 to 7:00 a.m. Sunday, July 5, 1998.

WEATHER: Partly cloudy. The temperature range was 59oF to 73oF.

I. CURRENT SITUATION: The contractor is working 24 hours a day, 7 days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 23 hours 46 minutes and 1006.2 tons of contaminated soils were treated during the time frame of this report. Site soil was processed at an average feed rate of 42.3 tons per hour. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: Continued TDF operations, feed preparation, debris separation, dust control, ash handling, and ash backfilling. Continued with real time perimeter air monitoring and off-site air monitoring activities. Received bottom ash analytical reports (bot-ash-06-29-98-151 and bot-ash-06-30-98-152) for bins 1 and 2 which met project specifications for backfill. Received fly ash analytical reports (fly-ash-06-29-98-026 from bin D which met project specifications for backfill above elevation 553.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF; continuation of excavation, drum handling, drum bulking, feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock washing & inorganic debris decontamination.

IV. PROBLEMS NOTED OR ANTICIPATED: None

V. NOTED VISITORS: None

PREPARED BY: Danny Zion for David Modricker (USACE).


Saturday, July 4, 1998

FOR DATES: From 7:00 a.m. Friday, July 3, 1998 to 7:00 a.m. Saturday, July 4, 1998.

WEATHER: Sunny and cool. The temperature range was 56oF to 82oF.

I. CURRENT SITUATION: The contractor is working 24 hours a day, 7 days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 23.8 hours and 1003.6 tons of contaminated soils was treated during the time frame of this report. Off-line time was attributed to the performance of the daily weight belt calibration. Site soil was processed at an average rate of 42.1 tons per hour. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: Continued TDF operations, feed preparation, debris separation, dust control, ash handling, and ash backfilling. Continued with real time perimeter air monitoring and off-site air monitoring activities.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF; continuation of excavation, drum handling, drum bulking, feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock washing & inorganic debris decontamination.

IV. PROBLEMS NOTED OR ANTICIPATED: None

V. NOTED VISITORS: None

PREPARED BY: Steve Gibson for David Modricker (USACE).


Friday, July 3, 1998

FOR DATES: From 7:00 a.m. Thursday, July 2, 1998 to 7:00 a.m. Friday, July 3, 1998.

WEATHER: Sunny and cool. The temperature range was 54oF to 80oF.

I. CURRENT SITUATION: The contractor is working 24 hours a day, 7 days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 23.75 hours and 1004.4 tons of contaminated soils were treated during the time frame of this report. Off-line time was attributed to performing the daily weight belt calibration. Site soil was processed at 42.3 tons per hour. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: Continued TDF operations, WWTP operations, excavation, feed preparation, debris separation, dust control, ash handling, ash and rock backfilling. Continued off-site and perimeter air monitoring activities. Received bottom ash analytical reports (bot-ash-06-28-98-150) for bin 7. Bottom ash tested met project specifications for backfill at or above elevation 547.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF; continuation of excavation, drum handling, drum bulking, feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock washing & inorganic debris decontamination.

IV. PROBLEMS NOTED OR ANTICIPATED: None

V. NOTED VISITORS: Charles Bragdon (USACE CIH)

PREPARED BY: John Ricard and David Modricker (USACE).


Thursday, July 2, 1998

FOR DATES: From 7:00 a.m. Wednesday, July 1, 1998 to 7:00 a.m. Thursday, July 2, 1998.

WEATHER: Sunny and breezy. The temperature range was 54oF to 79oF.

I. CURRENT SITUATION: The contractor is working 24 hours a day, 7 days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 23.5 hours and 993.8 tons of contaminated soil were treated during the time frame of this report. Off-line time was attributed to performing the daily weight belt calibration and relighting one of the Secondary Combustion Chamber burners. Site soil was processed at 42.3 tons per hour. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: Continued TDF operations, WWTP operations, excavation, feed preparation, debris separation, dust control, ash handling, ash backfilling and rock backfilling. Continued off-site and perimeter air monitoring activities. Received bottom ash analytical reports (bot-ash-06-26-98-148 and bot-ash-06-27-98-149) for bins 5 and 6. Bottom ash tested met project specifications for backfill at or above elevation 553.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF; continuation of excavation, drum handling, drum bulking, feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock washing & inorganic debris decontamination.

IV. PROBLEMS NOTED OR ANTICIPATED: None

V. NOTED VISITORS: Charles Bragdon (USACE CIH)

PREPARED BY: Steve Gibson and David Modricker (USACE).


Wednesday, July 1, 1998

FOR DATES: From 7:00 a.m. Tuesday, June 30, 1998 to 7:00 a.m. Wednesday, July 1, 1998.

WEATHER: Sunny with evening thunder storms. The temperature range was 61oF to 80oF.

I. CURRENT SITUATION: The contractor is working 24 hours a day, 7 days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 22.7 hours and 953.4 tons of contaminated soil were treated during the time frame of this report. Off-line time was attributed to performing the daily weight belt calibration and manually stopping feed in the evening due to a severe local thunderstorm. Site soil was being processed at 42 tons per hour. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: Continued TDF operations, WWTP operations, excavation, feed preparation, debris separation, ash handling, ash backfilling and rock backfilling. Continued off-site and perimeter air monitoring activities.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF; continuation of excavation, drum handling, drum bulking, feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock washing & inorganic debris decontamination.

IV. PROBLEMS NOTED OR ANTICIPATED: None

V. NOTED VISITORS: Sandy Downs, Larry Newcomer, Steve Webster, Mike Welch (PADEP), Christine Maggi, Rich Marcinkevage (City of Lock Haven).

PREPARED BY: Steve Gibson and David Modricker (USACE).

Region 3 | Mid-Atlantic Cleanup | Mid-Atlantic Superfund |EPA Home | EPA Superfund Homepage


Local Navigation


Jump to main content.